# Recommender systems and reinforcement learning for building control and occupant interaction: A text-mining driven review of scientific literature

### v1

@@ Abstract
The indoor environment significantly impacts human health and well-being; enhancing health and reducing energy consumption in these settings is a central research focus. With the advancement of Information and Communication Technology (ICT), recommendation systems and reinforcement learning have emerged as promising approaches to induce behavioral changes to improve the indoor environment and energy efficiency of buildings. This study aims to employ text-mining and Natural Language Processing (NLP) techniques to thoroughly examine the connections among these approaches in the context of building control and occupant interaction. The study analyzed around 27,000 articles from the ScienceDirect database, revealing extensive use of recommendation systems and reinforcement learning for space optimization, location recommendations, and personalized control suggestions. Although these systems are broadly applied to specific content, their use in optimizing indoor environments and energy efficiency remains limited. This gap likely arises from the need for interdisciplinary knowledge and extensive sensor data. Traditional recommendation algorithms, including collaborative filtering, content-based and knowledge-based methods, are commonly employed. However, the more complex challenges of optimizing indoor conditions and energy efficiency often depend on sophisticated machine learning (ML) techniques like reinforcement and deep learning. Furthermore, this review underscores the vast potential for expanding recommender systems and reinforcement learning applications in buildings and indoor environments. Fields ripe for innovation include predictive maintenance, building-related product recommendation, and optimization of environments tailored for specific needs, such as sleep and productivity enhancements based on user feedback. The study also notes the limitations of the method in capturing subtle academic nuances. Future improvements could involve integrating and fine-tuning pre-trained language models to better interpret complex texts.
@@
Reinforcement Learning , Recommender Systems , Building Operations , Building Energy Efficiency , Word Embeddings , Natural Language Processing
@@
Recommender systems and reinforcement learning for building control and occupant interaction: A text-mining driven review of scientific literature
@@
  • 27,000 full-text articles related to recommendation systems and the built environment have been extracted for text mining.
  • Detailed relationships among recommender algorithms, system types, input data, interventions, and platforms have been elucidated.
  • Word embeddings model has been used to extract the relationship between keywords.
  • The most commonly used algorithms, well-established applications, and emerging opportunities in building control and occupant interaction have been discussed.
@@ 1.1 Background
The indoor environment significantly affects human health and well-being, with individuals spending on average approximately 86% of their day indoors [1, 2]. Addressing how to improve health and well-being in indoor environments while reducing energy consumption is a central question in contemporary research [3]. In recent years, smart building control has become an important topic in this regard, with extensive studies demonstrating its ability to significantly reduce energy usage while maintaining indoor comfort levels [4, 5, 6, 7]. However, despite these promising results, the adoption rate of smart control technologies remains relatively low [8]. The primary barriers to widespread adoption are the high initial costs and complexity of implementation [9]. Therefore, there remains a need to explore a more feasible, user-friendly, and cost-effective solution.
@@
In the context of the widespread adoption of Artificial Intelligence (AI), the Internet of Things (IoT) devices, and smart mobile devices, a viable approach is the behavior change facilitated by Information and Communication Technology (ICT) [10, 11]. This method leverages recommendation systems and smart devices to influence and change human behavior, thus improving indoor environmental quality and reducing energy consumption [10]. Central to this strategy is sophisticated recommendation algorithms that analyze user preferences and environmental data to provide personalized suggestions. These algorithms have shown success in various sectors, including mobile health, commonly referred to as Just-In-Time Adaptive Interventions (JITAI) in the domain [12], as well as in online shopping [13], entertainment [14, 15], and social networking [16], etc.
@@
In the field of building and built environments, several conventional literature reviews have been conducted recently. These reviews analyze current studies on recommendation systems in various areas, including building performance optimization [17, 18, 19], IoT [20, 21, 22], and smart cities [23, 24, 25, 26]. These systems are typically utilized to optimize indoor environments [27, 28] and enhance building energy efficiency [29, 30, 31, 32]. Predominantly, the methods involve leveraging real-time data and federated learning to train models that generate energy-saving suggestions [33, 34, 35], alongside integrating sensor data with user preference feedback to provide personalized recommendations [36, 37]. Furthermore, reinforcement learning-based recommender systems are widely utilized in this domain, due to their ability to handle multi-objective optimization tasks [38, 39]. For example, multi-agent deep reinforcement learning has been developed to optimize energy consumption, occupant comfort, and air quality simultaneously in commercial buildings [40]. These systems are typically reactive, meaning they respond to user inputs or environmental sensors to provide suggestions. However, proactive recommendation systems have also been explored, which aim to incrementally alter user habits during specific moments based on the theories of micro-moments and habit loops [41]. Additionally, due to the opacity of machine learning (ML) black box models, explainable AI has been used to provide explainable and personalized energy efficiency suggestions [42]. On an urban scale, the application of recommendation systems has also garnered interest, primarily to improve energy efficiency in smart grids [43, 44] and facilitate demand response initiatives [45].
These reviews cover the diverse applications of recommendation systems in various aspects of the built environment. However, due to the limitations of human analytical capacity, conventional literature reviews may not adequately handle the vast volume of publications, struggling to discern the subtle relationships among different studies, particularly in interdisciplinary fields [46, 47]. Typically, these reviews are restricted to analyzing between 80-100 publications. Given the complexity and interdisciplinarity of integrating recommendation systems within the built environment, a more sophisticated approach is essential to efficiently manage and interpret extensive literature datasets. To address these challenges, text mining-based literature review methods emerge as a key tool.
@@ 1.2 Text mining-based literature review
Text mining-based literature reviews utilize data-driven technologies to systematically extract and analyze information from a large corpus of text, which is particularly advantageous in fields characterized by rapid technological advancements and extensive academic output [48]. Using natural language processing (NLP) and ML algorithms, text mining can efficiently process text data, including classification, clustering, and association, to identify trends, patterns, and emerging themes that might not be evident through traditional review methods [47, 49].
@@
Previous applications of text mining in literature reviews have demonstrated promising results. Tools such as VOSviewer [50] are commonly used in bibliometric analysis and network visualizations, using paper metadata to create bibliometric networks and density maps between articles in various fields [51, 52, 53]. However, these tools are often constrained by graphical user interfaces (GUIs), which limit the user’s ability to modify or extend beyond the provided algorithms [47]. To address specific text mining needs, researchers have adopted advanced ML techniques such as topic modeling. For example, Latent Dirichlet Allocation (LDA) [54] and the Correlated Topic Model (CTM) [55] have been effectively used to model documents as mixtures of topics [56, 57, 58]. Similarly, [59] employs the BERTopic model [60] to efficiently extract and cluster topics, thus providing deeper insights into textual data. Recent advances in computational power and algorithmic sophistication have also led to the emergence of Large Language Models (LLMs), such as GPT [61] and LLaMA [62], offering an alternative to traditional topic modeling approaches [63]. Furthermore, word embedding technologies such as Word2Vec [64, 65] and GloVe [66] facilitate character-level analysis [47]. These are complemented by tools such as Python and Scikit-Learn [67], RapidMiner Studio [68], N-gram [69], and the Natural Language Toolkit (NLTK) [70] for pre or post-processing, offering a nuanced understanding of word relationships and topic significance [71, 72].
@@
1.3 Review of recommender systems for the built environment using text-mining
@@
This study aims to overcome the challenges and limitations in typical literature reviews while comprehensively examining the interrelationships between recommendation systems in the context of building control and occupant interaction. Specifically, this study utilizes data mining techniques proposed by Abdelrahman et al. [47] to explore the relationships among five distinct categories of keywords. These categories include algorithms, types of recommendation systems, input data, interventions/objectives, and platforms, and Figure 1 illustrates the relational network between these categories. This represents the first effort to apply text mining and NLP to review these elements systematically in the context of the built environment and recommendation systems.
@@
Figure 1: Overview of the categories of concepts analyzed in this text-mining analysis and their relationships with each other.
@@ 2.1 Step 1: Article retrieval
ScienceDirect, operated by Elsevier, is a comprehensive academic database and full-text library that serves as major source for scholarly articles [73]. This study uses the Elsevier API [74] to perform full-text searches within the ScienceDirect database. The full text of each article and key metadata, including titles, abstracts, keywords, authors, and publication dates, are extracted and stored in JSON format. This format supports efficient data manipulation and retrieval for subsequent analysis phases.
@@ 2.1.1 Search query
The search strategy is designed to extract articles related to recommendation systems within built environments or related to indoor health. The query utilized combines specific terms to refine the focus of the retrieval: (recommendation system OR recommender system OR jitai) AND (energy efficiency OR indoor environment) AND (building OR built environment). The inclusion of terms (buiding OR built environment) ensures that the search results are relevant to the fields of building and built environment studies.
@@ 2.1.2 Article filtering
The initial query procedure yielded more than 60,000 articles sourced from more than 1,000 journals (the initial inquiry was conducted in May 2024). A significant proportion of these articles were found to be duplicated and some contained incomplete or erroneous information. Following the removal of these duplicates and erroneous files, the dataset was refined to approximately 27,000 articles. Each article retained includes the full text along with essential metadata such as the author(s), publication year, title, abstract, and journal of publication. Figure 3 displays the distribution of the number of articles per journal and the trends of annual publication. Analysis of these data reveals that most of the articles originate from journals focused on the construction, energy, and environment sectors. In addition, a smaller subset of these articles is derived from medical journals.
@@ 2.2.1 Definitions
    Recommender system: (recommende_system) refers to the types of recommender systems commonly used in the field of the built environment. This encompasses systems like JITAI, context-aware recommender systems for dynamic user environments, and reinforcement learning-based recommender systems that adapt based on user interactions.
@@
    Algorithm: (algorithm) refers to the types of algorithms employed across various recommender systems. Notable examples include NLP, deep learning, and reinforcement learning.
@@
    Objective: (objective) refers to the goals pursued by recommender systems under the context of the built environment. Examples include promoting energy-saving behavior, altering sedentary behaviors, or providing personalized control recommendations.
@@ 2.3 Step 3: Text pre-possessing
The pre-processing stage is designed to refine the full-text data for subsequent data mining processes. The preparation involves removing words that could negatively impact model training, merging compound words, and performing word tokenization and lemmatization. Initially, two types of undesirable words are identified: 1) metadata elements such as article or image IDs, author information, hyperlinks, and annotations, including terms such as introduction, methodology, reference, figure, and table, which are repetitive between articles and can introduce bias in subsequent analyses; 2) stop words, which occur frequently but contextually insignificant words such as the, a, in, and of, which if included, could bias the NLP models. NLTK is then used to remove these stop words [70].

### v2

# Recommender systems and reinforcement learning for human-building interaction and context aware support: A text mining-driven review of scientific literature

@@ Abstract
The indoor environment significantly impacts human health and well-being; enhancing health and reducing energy consumption in these settings is a central research focus. With the advancement of Information and Communication Technology (ICT), recommendation systems and reinforcement learning (RL) have emerged as promising approaches to induce behavioral changes to improve the indoor environment and energy efficiency of buildings. This study aims to employ text mining and Natural Language Processing (NLP) techniques to thoroughly examine the connections among these approaches in the context of human-building interaction and occupant context-aware support. The study analyzed 27,595 articles from the ScienceDirect database, revealing extensive use of recommendation systems and RL for space optimization, location recommendations, and personalized control suggestions. Although these systems are broadly applied to specific content, their use in optimizing indoor environments and energy efficiency remains limited. This gap likely arises from the need for interdisciplinary knowledge and extensive sensor data. Traditional recommendation algorithms, including collaborative filtering, content-based and knowledge-based methods, are commonly employed. However, the more complex challenges of optimizing indoor conditions and energy efficiency often depend on sophisticated machine learning (ML) techniques like reinforcement and deep learning. Furthermore, this review underscores the vast potential for expanding recommender systems and RL applications in buildings and indoor environments. Fields ripe for innovation include predictive maintenance, building-related product recommendation, and optimization of environments tailored for specific needs, such as sleep and productivity enhancements based on user feedback. The study also notes the limitations of the method in capturing subtle academic nuances. Future improvements could involve integrating and fine-tuning pre-trained language models to better interpret complex texts.
@@
Reinforcement Learning , Recommendation Systems , Human-Building Interaction , Occupant-Centric , Building Energy Efficiency , Word Embeddings , Natural Language Processing
@@
Recommender systems and reinforcement learning for human-building interaction and context-aware support: A text mining-driven review of scientific literature
@@
  • 27,595 full-text articles related to the application of recommender systems in buildings and occupant interactions have been extracted for text mining.
  • Detailed relationships between recommender algorithms, system types, input data, interventions, and platforms have been elucidated.
  • Word embeddings model has extracted the relationship between keywords.
  • The most widely used algorithms, well-established applications, and emerging opportunities in human-building interaction and occupant context-aware support have been discussed.
@@ 1.1 Background
The indoor environment significantly affects human health and well-being, with individuals spending on average approximately 86% of their day indoors [1, 2]. The COVID-19 pandemic has further highlighted the critical importance of indoor air quality, propelling research focused on mitigating airborne transmission of pathogens within buildings [3]. For example, [4] has proposed a tailored approach to calculate the optimal number of outdoor air changes in HVAC systems for school buildings to address post-pandemic challenges. However, some studies have also highlighted that improved ventilation systems, while enhancing occupant health, can impact energy consumption and operational costs [5, 6]. Consequently, a critical question in contemporary research is how to balance the dual objectives of improving health and well-being in indoor environments while minimizing energy consumption [7]. In recent years, smart building control has become an important topic in this regard, with extensive studies demonstrating its ability to significantly reduce energy usage while maintaining indoor comfort levels [8, 9]. However, despite these promising results, the adoption rate of smart control technologies remains relatively low [10]. The primary barriers to widespread adoption are the high initial costs and complexity of implementation [11]. Therefore, there remains a need to explore a more feasible, user-friendly and cost-effective solution.
@@
In the context of the widespread adoption of Artificial Intelligence (AI), the Internet of Things (IoT) devices, and smart mobile devices, a viable approach is the behavior change facilitated by Information and Communication Technology (ICT) [12, 13]. This method leverages recommendation systems and smart devices to influence and change human behavior, thus improving indoor environmental quality and reducing energy consumption [12]. Central to this strategy are sophisticated recommendation algorithms that analyze user preferences and environmental data to provide personalized suggestions. These algorithms have shown success in various sectors, including mobile health, commonly referred to as Just-In-Time Adaptive Interventions (JITAI) in the domain [14], as well as in online shopping [15], entertainment [16, 17], and social networking [18], etc.
@@
In the context of building and human-building interactions, these systems are typically utilized to optimize indoor environments [19, 20] and enhance building energy efficiency [21, 22, 23]. Mostly, the methods involve the use of real-time data and federated learning to train models that generate energy saving suggestions [24, 25, 26], alongside the integration of sensor data with user preference feedback to provide personalized recommendations [27, 28]. Furthermore, reinforcement learning-based recommender systems are widely utilized in this domain due to their ability to handle multi-objective optimization tasks [29]. For example, multi-agent deep reinforcement learning (RL) has been developed to optimize energy consumption, occupant comfort, and air quality simultaneously in commercial buildings [30]. These systems are typically reactive, meaning they respond to user inputs or environmental sensors to provide suggestions. However, proactive recommendation systems have also been explored, which aim to incrementally alter user habits during specific moments based on the theories of micro-moments and habit loops [31]. Furthermore, due to the opacity of machine learning (ML) models black box, explainable AI has been used to provide explainable and personalized energy efficiency suggestions [32]. On an urban scale, the application of recommendation systems has also garnered interest, primarily to improve energy efficiency in smart grids [33, 34] and facilitate demand response initiatives [35].
Ref. Year Sector Focus Contributions
[36] 2021 Building 1) algorithm types; 2) algorithm design; 3) applications; 4) limitations; 5) future directions 1) The first critical review of energy efficiency recommendation systems in buildings; 2) Provides classification and identifies current challenges and unresolved issues, and offers insights into future directions.
[37] 2021 Energy services 1) algorithm types; 2) platforms; 3) user profiles; 4) applications 1) Reviews recommendation system’s applications in demand-side management, smart services, e-commerce, residential, industrial, commercial, and policy contexts.
[38] 2019 IoT 1) literature quantification analysis; 2) system design; 3) limitations; 4) future directions 1) Examines recommendation techniques in IoT environments; 2) Highlights advantages, disadvantages, and unresolved challenges for various methods.
[39] 2021 IoT 1) literature quantification analysis; 2) algorithm types 3) applications; 4) algorithm design; 5) future directions 1) Comprehensively reviews IoT-based recommendation technologies; 2) Summarizes applications in various fields and identifies challenges; 3) Provides an RSIoT reference framework to guide future research and practices.
[40] 2020 Smart cities 1) algorithm types; 2) system design; 3) applications; 4) literature quantification analysis 1) Reviews the status and key drivers of recommendation systems for smart cities across domains like economy, environment, mobility, and living.
[41] 2023 Smart cities 1) literature quantification analysis; 2) applications; 3) future directions 1) Highlights recommendation systems in energy efficiency, government services, healthy living, parking optimization, and traffic congestion reduction.
[42] 2024 Smart cities 1) literature quantification analysis; 2) algorithm types; 3) applications 1) Explores trends, themes, and collaborations in smart city recommendation systems; 2) Identifies cooperation between authors and institutions.
[43] 2024 Smart cities 1) literature quantification analysis; 2) algorithm types; 3) applications 1) Summarizes recommendation systems in tourism, health, mobility, and transport; 2) Highlights common algorithms and application scenarios.
Table 1: Summary of Literature on Recommendation Systems
1.2 Previous reviews and research gaps
Recommendation systems have been widely applied to improve various aspects of building performance and energy management and several literature reviews have recently been conducted to summarize the applications of recommendation systems in the context of building and the built environment. These reviews analyze current studies on recommendation systems in various areas, including energy efficiency [36, 37], Internet of Things (IoT) [38, 39], and smart cities [40, 41, 42, 43]. Table 1 summarizes the focus areas and key contributions of these review articles. From this table, it is evident that most reviews analyze the application of recommendation systems in the field of building and the built environment within the context of smart cities and the IoT, and only one study has reviewed their application at the building scale. However, this study exclusively focuses on the use of recommendation systems for building energy efficiency [12]. It does not explore their potential contributions to other core objectives of building-human interaction design, such as improving occupant wellness and comfort. Furthermore, these reviews focus primarily on analyzing the individual components of recommendation systems (typically data sources, algorithm types, system types, platforms, and recommendation objectives) in isolation. They fail to investigate the interrelationships among these categories, such as the connections between different objectives and data types or algorithms, which is crucial for guiding the design of recommendation systems tailored to specific objectives. This issue often arises due to the limitations of human analytical capacity, as it is challenging to discern subtle relationships between different studies, particularly in interdisciplinary fields [44, 45]. Therefore, to address this gap and investigate the interrelations between various components, a more sophisticated approach is essential to efficiently manage and interpret extensive literature datasets.
@@ 1.3 Text mining-based literature review
Text mining-based literature reviews utilize data-driven technologies to systematically extract and analyze information from a large corpus of text, which is particularly advantageous in fields characterized by rapid technological advancements and extensive academic output [46]. Using natural language processing (NLP) and ML algorithms, text mining can efficiently process text data, including classification, clustering, and association, to identify trends, patterns, and emerging themes that might not be evident through traditional review methods [47].
@@
Previous applications of text mining in literature reviews have demonstrated promising results. Tools such as VOSviewer [48] are commonly used in bibliometric analysis and network visualization [49], using paper metadata to create bibliometric networks and density maps between articles in various fields [50, 51]. However, these tools are often constrained by graphical user interfaces (GUIs), which limit the user’s ability to modify or extend beyond the provided algorithms [45]. To address specific text mining needs, researchers have adopted advanced ML techniques such as topic modeling. For example, Latent Dirichlet Allocation (LDA) [52] and the Correlated Topic Model (CTM) [53] have been effectively used to model documents as mixtures of topics [54, 55, 56]. Similarly, [57] employs the BERTopic model [58] to efficiently extract and cluster topics, thus providing deeper insights into textual data. Recent advances in computational power and algorithmic sophistication have also led to the emergence of Large Language Models (LLMs), such as GPT [59] and LLaMA [60], offering an alternative to traditional topic modeling approaches [61]. Furthermore, word embedding technologies such as Word2Vec [62] and GloVe [63] facilitate character-level analysis [45]. These are complemented by tools such as Python and Scikit-Learn [64], RapidMiner Studio [65], N-gram [66], and the Natural Language Toolkit (NLTK) [67] for pre or post-processing, offering a nuanced understanding of word relationships and topic significance [68, 69].
Among these methods, Word2Vec is frequently used by researchers to quantify the relationships between terms associated with different topics from large volumes of literature. For instance, [70] developed an unsupervised embedding model based on Word2Vec using abstracts from 3.3 million papers to extract materials science information and predict new thermoelectric materials; [71] utilized abstracts from nearly 4 million publications as training data to explore the relationships between cancer types and anticancer drugs; and [72] employed 4,633 articles retrieved from the Web of Science database to develop a Word2Vec model, uncovering latent patterns and interactions within resources, conservation and recycling research topics. In the context of building energy efficiency, [45] retrieved approximately 30,000 scientific publications via the Elsevier API to extract relationships among data sources, data science techniques, and applications for building energy efficiency across the entire lifecycle of buildings. This study represents the first attempt at such a text mining and NLP-based review in the field of building energy efficiency research.
@@
1.4 Review of recommender systems for human-building interaction using text mining
@@
This study aims to leverage data mining techniques proposed by Abdelrahman et al. [45] to comprehensively review the applications of recommendation systems in the context of human-building interaction and occupant context-aware support. Specifically, it seeks to quantify the interrelationships among five key components of recommendation systems: algorithms, types of recommendation systems, input data, interventions/objectives, and platforms. Figure 1 illustrates six specific relationships among these five categories, which have been identified through prior literature reviews.
@@
Figure 1: Overview of the categories of concepts analyzed in this text mining analysis and their relationships with each other.
The objectives of this study are as follows: 1) to train a Word2Vec model on relevant academic papers to extract keyword relationships; 2) to create heatmaps that visualize the interrelationships between keywords within the different components of recommendation systems; 3) to analyze the strengths and weaknesses of these relationships as shown in the heatmaps, thereby identifying the current state of development in the field. The main contributions of this study can be summarized as follows:
  1. 1.
    This study is the first to provide a comprehensive review of recommendation systems in the context of human-building interaction and occupant context-aware support.
  2. 2.
    Quantifies the correlations between recommendation objectives, input data, algorithms, and platforms by using NLP and text mining techniques.
  3. 3.
    The findings offer practical guidance on selecting appropriate data and algorithms to design recommendation systems that target different objectives.
  4. 4.
    Explores the most widely used recommendation algorithms, application areas, and future development directions.
@@ 2.1 Step 1: Article retrieval
ScienceDirect, operated by Elsevier, is a comprehensive academic database and full-text library that serves as the main source for scholarly articles [73]. This study uses the Elsevier API [74] to perform full-text searches within the ScienceDirect database. The full text of each article and key metadata, including titles, abstracts, keywords, authors, and publication dates, are extracted and stored in JSON format. This format supports efficient data manipulation and retrieval for subsequent analysis phases.
@@ 2.1.1 Search query
The search strategy is designed to extract articles related to the application of recommender systems in buildings and occupant interactions. The query utilized combines specific terms to refine the focus of the retrieval: (recommendation system OR recommender system OR jitai) AND (energy efficiency OR indoor environment) AND (building OR built environment). The inclusion of terms (building OR built environment) ensures that the search results are relevant to the fields of building and built environment studies.
@@ 2.1.2 Article filtering
The initial query procedure yielded more than 60,000 articles sourced from more than 1,000 journals (the initial inquiry was conducted in May 2024). A significant proportion of these articles were found to be duplicated and some contained incomplete or erroneous information. Following the removal of these duplicates and erroneous files, the data set was refined to 27,595 articles. Each retained article includes the full text along with essential metadata such as the author(s), publication year, title, abstract, and journal of publication. Figure 3 displays the distribution of the number of articles per journal and the trends of annual publication. Analysis of these data reveals that most of the articles originate from journals focused on the construction, energy, and environment sectors. In addition, a smaller subset of these articles is derived from medical journals.
@@ 2.2.1 Definitions
    Recommender system: (recommende_system) refers to the types of recommender systems commonly used in the field of human-building interaction. This encompasses systems like JITAI, context-aware recommender systems for dynamic user environments, and RL-based recommender systems that adapt based on user interactions.
@@
    Algorithm: (algorithm) refers to the types of algorithms employed across various recommender systems. Notable examples include NLP, deep learning, and RL.
@@
    Objective: (objective) refers to the goals pursued by recommender systems under the context of human-building interaction. Examples include promoting energy-saving behavior, altering sedentary behaviors, or providing personalized control recommendations.
@@ 2.3 Step 3: Text pre-possessing
The pre-processing stage is designed to refine the full-text data for subsequent data mining processes. The preparation involves removing words that could negatively impact model training, merging compound words, and performing word tokenization and lemmatization. Initially, two types of undesirable words are identified: 1) metadata elements such as article or image IDs, author information, hyperlinks, and annotations, including terms such as introduction, methodology, reference, figure, and table, which are repetitive between articles and can introduce bias in subsequent analyses; 2) stop words, which occur frequently but contextually insignificant words such as the, a, in, and of, which, if included, could bias the NLP models. NLTK is then used to remove these stop words [67].